Ball A, with a mass of 1000 g moving in a constant velocity of 30 m/s, hits Ball B with a mass of 4000 g which is at rest. As a result of the collision, Ball A bounced back at 22 m/s. determine the final velocity of Ball B.



1
Expert's answer
2021-12-13T10:34:29-0500

By momentum conservation:


"m_Av_A+m_B·0=-m_Au_A+m_Bu_B,\\\space\\
u_B=\frac{m_A(v_A+u_A)}{m_B}=13\text{ m/s}."